OpenStudy (anonymous):

If the volume of wood is 15.84 cubic feet when the height is 22 ft and the distance around is 3ft, what is the volume of wood obtained from a tree that is 37 ft tall and having a measurement of 4ft around the trunk?

OpenStudy (anonymous):

is it a cylinder?

OpenStudy (anonymous):

yeah

OpenStudy (campbell_st):

find the radius.... using the circumference C = 2pi r 4 = 2pi r r= 2/pi feet Volume: \[V = \pi \times (2/\pi)^2 \times 37\] evaluate the volume for your answer